OpenStudy (anonymous):

Find the sum of the geometric sequence. 1, one divided by four, one divided by sixteen, one divided by sixty four, one divided by two hundred and fifty six

OpenStudy (anonymous):

1/4 +1/16 +1/64 these terms are in 4's power

OpenStudy (anonymous):

can you tell me whats the common ratio

OpenStudy (anonymous):

honestly i dont know

OpenStudy (anonymous):

to find the common ratio, take the second term and divide it by the first term. (1/16) / (1/4 = 1/16 * 4/1 = 4/16 = 1/4 so the common ratio is 1/4. Is this all the numbers ? Is there only 4 numbers that are trying to find the sum of ?

OpenStudy (anonymous):

there 5 number 1, 1/4 ,1/16, 1/64 , 1/256

OpenStudy (anonymous):

oh...1 is the first term ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so

OpenStudy (anonymous):

yes still the same ratio

OpenStudy (anonymous):

so what is the the sum of the geometric sequence.

OpenStudy (anonymous):

you cant make bricks without clay,you cant make buildings without bricks data data! how can you find something without data

OpenStudy (anonymous):

you have the ratio now now formula is a(1-r^n/1-r)

OpenStudy (anonymous):

first term is 1 r= 1/4

OpenStudy (anonymous):

and n is 5

OpenStudy (anonymous):

so sum will be (1-1/256)/1-1/4

OpenStudy (anonymous):

so 255/4 *4/3 = 255/3 = 85

OpenStudy (anonymous):

thats ur sum

OpenStudy (anonymous):

341/256 is the answer. I just took the test and got it right.